The radius of a silver bar is 30 π‘π‘š. What force is required to stretch the bar by 12.7% of its length where Young’s modulus is π‘Œ = 7.25 Γ— 1010 π‘π‘šβˆ’2 .


1
Expert's answer
2021-12-06T09:47:35-0500

The strain of 12.7%, or 0.127, is what we will use to solve the problem:


"\epsilon=\frac{\sigma}{E}=\frac{F}{AE},\\\space\\
F=\epsilon AE=\epsilon·\pi r^2·E,\\\space\\
F=2.6\text{ MN}."